# A Control Allocation Algorithm for Hypersonic Glide Vehicles with Input Limitations

### v2

@@ 2.1.2 Redistributed Pseudoinverse-Based Control Allocation
An alternative approach to solve the constrained control allocation problem was proposed by the authors in [20] and is called redistributed pseudoinverse-based control allocation (RPICA). The core idea of the methodology is to cope with the problem of limited control input authority by iterating over the AMS until a feasible control input command is found. The algorithm starts by generating a first solution for a control vector candidate using the pseudoinverse
@@
with $\bm{\nu}^{(0)}$ being the initially desired commanded virtual control input vector from the controller. In the case where no element of $\mathbf{u}^{(1)}$ is saturated, and Eq. (7) delivers the desired virtual control input vector, $\mathbf{u}^{(1)}$ is the final solution. In cases where at least one element of $\mathbf{u}^{(1)}$ conflicts with the available control limitations and the desired virtual control input vector is not generated, the process continues. The achievable control input is defined as
@@
For MIMO systems, it is preferable to keep the directionality of the unconstrained input vector $\mathbf{u}^{(j)}$. To make that possible, the authors in [21] proposed the redistributed scaling pseudoinverse control allocation (RSPICA) algorithm. The saturated control input vector is scaled within the method to ensure perfect direction alignment with the desired pseudo-control vector. The used scalar $a$ is chosen such that the element that exceeds its respective limit the most is saturated, while the others are not:
@@
Both RPICA and its scaled variant, RSPICA, are advantageous for implementation in embedded systems with limited computational resources due to their simplicity. RSPICA, in particular, was designed to maintain input vector directionality in MIMO systems under saturation conditions by scaling the requested vector with a factor $a$, thereby ensuring that control input constraints are respected. However, despite these advancements, both methods have notable limitations. Their reliance on the $\ell_{2}$-based least-squares approach, inherent in the Moore-Penrose pseudoinverse, can result in an inefficient use of the AMS, when the asymmetrical input constraints and strong inter-directional control authority of the control effectors are present. Additionally, for systems with asymmetric input limits where the lower input bound is zero, RSPICA may produce zero-scaled solutions, resulting in trivial control commands. These challenges are particularly problematic for the here considered class of HGVs.
@@ 2.1.3 Quadratic Programming-Based Control Allocation
The constrained control allocation problem can also be solved using a quadratic programming-based control allocation (QPCA) approach. Examples of that can be found in the work presented in [4] and [22]. To discuss this, we introduce the general formulation of a quadratic programming problem
@@
with $\mathbf{A}_{u}=[\mathbf{I}_{m\times m},-\mathbf{I}_{m\times m}]^{T}$ and $\mathbf{b}_{u}=[\min(u_{\max,i},\overline{u}_{i}),-\max(u_{\min,i},\overline{u}_{i})]^{T}$. The formulated QPCA problem can be solved using a variety of methods, such as interior-point, active-set, or simplex [5]. The advantage of the method is that it is able to provide solutions to complex control allocation problems. However, a significant drawback of the approach is the high computational complexity and with that the lack of real-time guarantees for certain embedded systems integrated in aerial vehicles, for example the system integrated into the GHGV-2.
@@ 3 Problem Formulation
Figure 5: Rear view of the GHGV-2 showing available control effectors during endoatmospheric operations [23].
Figure 5 illustrates the available aerodynamic control surfaces and their corresponding deflections: upper left fin $u_{1}$, upper right fin $u_{2}$, lower left fin $u_{3}$, and lower right fin $u_{4}$. This leads to the following control input vector
@@
with $\mathbf{B}(\bm{x}(t),\bm{\sigma}(t))\in\mathbb{R}^{3\times 4}$. As it is a non-square matrix, it cannot be inverted directly.
A major challenge in hypersonic flights, often neglected in the literature, is the dependence of actuator magnitude and rate limits not only on the instantaneous flight state $\bm{x}(t)$ but also on the operating point $\bm{\sigma}(t)$. During high-speed atmospheric operations, high dynamic pressure induces counterforces on actuators, limiting the achievable deflections and rates. Nonlinear aerodynamic effects further complicate matters, with asymmetric influences on upper and lower surfaces requiring distinct actuator constraints. Additionally, operating-point-dependent thermal loads must be considered to minimize heating, reduce the burden on the TPS, and lower the infrared signature of the vehicle during operations.
@@
These effects are illustrated in Fig. 6. In this figure, the red arrow $\mathbf{V}$ represents the resulting velocity vector of the vehicle as it moves at high speed through the atmosphere. The blue arrows depict the airflow vector fields around the vehicle, shaping the aerodynamic forces and heating effects on the control surfaces. The control surface $i$, modeled as a flapping plate, is characterized by a deflection angle $u_{i}$ and a deflection rate $\dot{u}_{i}$. The red gradient along the control surface indicates aerodynamic heating, generated by the vehicle’s hypersonic speed, causing significant heat flux $\tau_{i}$ on the surface. These heating effects must be considered during operation to avoid material degradation and structural failure of the control surfaces and to minimize the infrared signature in defense scenarios. The green arrow $F_{q,i}$ represents the normal of the aerodynamic force acting on the surface, which is highly dependent on the deflection angle and the dynamic pressure $q=0.5\rho V^{2}$ acting on the control surface area $S_{u,i}$, with $\rho$ being the atmospheric density and $V$ the resulting velocity. The actuator force $\mathbf{F}_{A,i}$ is responsible for controlling the position of the surface, working against $\mathbf{F}_{q,i}$, especially under high dynamic pressure conditions. Under these conditions, the actuator’s ability to adjust the deflection angle and rate is sometimes restricted, leading to cases in which the deflection angle $u_{i}$ is limited such that the resulting dynamic pressure force $F_{q,i}$ exceeds the maximum force that the actuator can provide.
@@
with $\bm{u}_{s}(t)$ being a desired steady-state solution for a given virtual control input $\bm{\nu}(t)$. $\mathbf{W}_{m}$ is a weighting matrix for the difference between the computed input and the desired steady-state solution, and $\mathbf{W}_{r}$ is a weighting matrix for the delta between the computed control input $\bm{u}(t)$ and the control input of the previous time step $\bm{u}(t-T)$.
In summary, the problem addressed in this work is to find, at each time, a control input $\bm{u}(t)$ for the system defined in Eqs. (1) and (2), in accordance with Eq. (31), such that the virtual control input demand $\mathbf{B}(\bm{x}(t),\bm{\sigma}(t))\,\bm{u}(t)=\bm{\nu}(t)$ is met if the command is within the state- and operating-point-dependent actuator limits. If the command is not feasible, the system should generate a command that minimizes the error between the requested and achievable command while respecting the input limitations from Eq. (30).
@@ 4 The Integrated Overall Control Allocation Concept
As discussed in Sect. 3, during atmospheric flight, HGVs face significant challenges in maintaining stability and control within a narrow altitude band due to either thermal loads or a lack of aerodynamic controllability [24]. These conditions complicate the design of effective control strategies and demand precise management of aerodynamic and thermal effects. Owing to this complexity, the atmospheric flight trajectory of the HGV is typically computed offline prior to the mission. During this offline optimization process, an optimal control problem is solved to ensure that, given specific initial and terminal conditions as well as physical constraints, a feasible trajectory is obtained [25].
@@
The resulting optimized state reference trajectory is bounded by aerodynamic flight corridor constraints, as illustrated in Fig. 7. This figure provides an example of an operational flight corridor for a hypersonic glide vehicle and its associated reference trajectory. The corridor is defined by upper and lower Mach number limits as a function of altitude. When the vehicle ascends above the upper altitude bound, it enters a flight region with insufficient aerodynamic control authority due to reduced air density. Conversely, descending below the lower bound places the vehicle in a region with excessive heating, risking material degradation and structural failure. Based on the obtained state reference trajectory $\bm{x}_{r}(t)$, an offline solution for the control input $\bm{u}_{r}(t)$ is computed using a computationally demanding but accurate QP solver for the problem defined in Eq. (31), employing the methods discussed in Sect. 2.1.3. This offline solution incorporates all known hard and soft actuator constraints, producing what is assumed to be the best solution for a disturbance-free environment. The computed control inputs from this offline optimization are stored in lookup tables and serve as baseline control inputs during flight.
The baseline control input $\bm{u}_{r}(t)$ would be sufficient to maintain the vehicle on its reference trajectory in a scenario without external disturbances or model uncertainties. However, given the inherent uncertainties and potential disturbances in hypersonic flight, the control system also requires an attitude feedback controller to provide corrective actions [23]. This controller computes a corrective virtual control input $\Delta\bm{\nu}(t)$ to compensate for deviations from the nominal trajectory due to unmodeled dynamics, uncertainties, and external disturbances. The proposed iterative online algorithm uses the virtual control input $\Delta\bm{\nu}(t)$ to search in the attainable moment set (AMS) for a feasible control input $\Delta\bm{u}(t)$. Assuming an affine structure for the control input, the final command can be expressed as a superposition of the offline baseline solution and the correction:
@@
The overall control allocation strategy is illustrated in Fig. 8. This structure enables the vehicle to respond dynamically to uncertainties while maintaining stability and control along its planned trajectory.
@@
In summary, combining an offline-optimized control input set with an online feedback adjustment ensures that the vehicle operates within its designated flight envelope while adapting in real time to maintain performance under challenging conditions. The corrective control input $\Delta\bm{u}(t)$, requested by the attitude feedback controller, is determined through an online optimization algorithm. The design and implementation of this online optimization algorithm are central contributions of this paper, and further details on its development and operation are presented in the following section.
@@ 5 The Proposed Online Control Allocation Algorithm
In standard redistributed pseudoinverse control allocation (RPICA) approaches, the problem of limited control input authority is solved by iterating over the admissible maneuvering set (AMS) until a feasible control input command is found [20]. To do so, the algorithm typically uses the Moore–Penrose inverse. The benefit of this approach is that a solution can often be obtained in a simple way that is real-time feasible for many aerospace applications. However, in most RPICA formulations no soft constraints on the distribution of the control effort are considered, and asymmetric input or rate limits cannot be represented adequately. For the class of HGVs studied here, these aspects are particularly critical, rendering standard RPICA solutions suboptimal.
In the proposed iterative dynamic control allocation (IDCA) algorithm, a linear filter formulation, originally introduced in [22], is employed. Unlike the Moore–Penrose inverse, this approach explicitly accounts for both the current control distribution and the allocation of the previous time step as soft constraints [26]. This implicitly incorporates actuator rates into the allocation problem. The method provides an explicit solution to the following weighted least-squares problem:
@@
where $\bm{\nu}(t)=\mathbf{B}\,\bm{u}_{r}(t)+\Delta\bm{\nu}(t)$ is the virtual control input to be generated. The desired steady-state solution $\bm{u}_{s}(t)$ is specified in Sect. 5.2. The algorithm is designed so that, if $\bm{u}_{s}(t)$ is feasible, the steady-state solution converges to $\bm{u}_{s}(t)$ or to a solution close to it. The weighting matrices $\mathbf{W}_{m}$ and $\mathbf{W}_{r}$ are symmetric and defined as functions of the state and control; their detailed design is discussed in Sect. 5.1.
Since $\mathbf{W}_{m}$ and $\mathbf{W}_{r}$ are symmetric, the combined weighting
@@
The closed-form solution of Eq. (33), expressed in the form of Eq. (35), can be written as [22]:
@@
The explicit solution of Eq. (36) is used within an iterative procedure to enforce actuator constraints. The algorithm is executed within a fixed controller cycle, so external variables such as $\bm{u}(t-T)$ and $\bm{x}(t)$ are treated as constant during the inner iterations. A closely related line of work was presented by Burken et al. [6], who introduced a fixed-point, quadratic-programming-based iteration scheme for control allocation. Their method shares structural similarities with the iterative procedures considered here, in that both refine the solution by re-evaluating residual commands under actuator constraints. While the fixed-point scheme demonstrated promising performance for asymmetric re-entry vehicle configurations and provided a rigorous QP-based foundation for handling magnitude constraints, actuator rate limits were only mentioned as a possible extension and not explicitly included. By contrast, the algorithm presented in the following integrates rate-dependent limits directly into the allocation problem, both as explicit constraints and as soft penalties in the weighted least-squares cost. Moreover, the proposed formulation emphasizes practical implementation by illustrating how the weighting structure can incorporate operational considerations such as thermal load reduction and drag-sensitive allocation, and by providing empirical evidence in the simulation results that the method can robustly handle strongly time-varying, state- and operating-point-dependent actuator constraints.
@@
We define the iterative residual $\bm{\nu}^{(j)}$. Initially, the control input $\bm{u}^{(1)}$, corresponding to $\bm{\nu}^{(0)}=\bm{\nu}(t)$, is computed from Eq. (36). If $\bm{u}^{(1)}$ respects all actuator constraints and delivers the required virtual control, it becomes the final solution. Otherwise, the process continues.
The achieved saturated virtual control in iteration step $j$ is
@@
where $S(\cdot)$ is the state- and rate-dependent saturation function (cf. Eq. (29)), $\dot{\bm{u}}^{(j)}=\tfrac{1}{T}(\bm{u}^{(j)}-\mathbf{u}_{\tau})$ and $\mathbf{u}_{\tau}=\mathbf{u}(t-T))$ die control input of the last time step.
If actuators saturate, the corresponding columns of $\mathbf{B}$ are set to zero, yielding the reduced matrix $\mathbf{B}_{s}$. Input limits are updated at each step to ensure feasibility:
@@
For completeness, alongside the mathematical steps and the process visualization in Fig. 9, the algorithm is also summarized in pseudocode to complement the formulation and facilitate implementation.

### v3

@@ 2.2.2 Redistributed Pseudoinverse-Based Control Allocation
An alternative approach to solve the constrained control allocation problem was proposed by the authors in [Virning_1994] and is called redistributed pseudoinverse-based control allocation (RPICA). The core idea of the methodology is to cope with the problem of limited control input authority by iterating over the AMS until a feasible control input command is found. The algorithm starts by generating a first solution for a control vector candidate using the pseudoinverse
@@
with $\bm{\nu}^{(0)}$ being the initially desired commanded virtual control input vector from the controller. In the case where no elements of $\mathbf{u}^{(1)}$ are saturated and Eq. (13) delivers the desired virtual control input vector $\nu^{(0)}$, $\mathbf{u}^{(1)}$ is the final solution. In cases where at least one element of $\mathbf{u}^{(1)}$ conflicts with the available control limitations and the desired virtual control input vector is not generated, the process continues. The achievable control input is defined as
@@
For MIMO systems, it is preferable to keep the directionality of the unconstrained input vector $\mathbf{u}^{(j)}$. To make that possible, the authors in [Zhang_2018] proposed the redistributed scaling pseudoinverse control allocation (RSPICA) algorithm. The saturated control input vector is scaled within the method to ensure perfect direction alignment with the desired pseudo-control vector. The used scalar $a$ is chosen such that the element that exceeds its respective limit the most is saturated, while the others are not:
@@
Both RPICA and its scaled variant, RSPICA, are advantageous for implementation in embedded systems with limited computational resources due to their simplicity. RSPICA, in particular, was designed to maintain input vector directionality in MIMO systems under saturation conditions by scaling the requested vector with a factor $a$, thereby ensuring that control input constraints are respected. However, despite these advancements, both methods have notable limitations. Their reliance on the $\ell_{2}$-based least-squares approach, inherent in the Moore-Penrose pseudoinverse, can result in an inefficient use of the AMS, when the asymmetrical input constraints and strong inter-directional control authority of the control effectors are present. Additionally, for systems with asymmetric input limits where the lower input bound is zero, RSPICA may produce zero-scaled solutions, resulting in trivial control commands. These challenges are particularly problematic for the certain classes of hypersonic vehicle.
@@ 2.2.3 Quadratic Programming-Based Control Allocation
The constrained control allocation problem can also be solved using a quadratic programming-based control allocation (QPCA) approach. Examples of that can be found in the work presented in [Johansen_2004] and [Ola_2004]. To discuss this, we introduce the general formulation of a quadratic programming problem
@@
with $\mathbf{A}_{u}=[\mathbf{I}_{m\times m},-\mathbf{I}_{m\times m}]^{T}$ and $\mathbf{b}_{u}=[\min(u_{\max,i},\overline{u}_{i}),-\max(u_{\min,i},\overline{u}_{i})]^{T}$. The formulated QPCA problem can be solved using a variety of methods, such as interior-point, active-set, or simplex [Ola2004]. The advantage of the method is that it is able to provide solutions to complex control allocation problems. However, a significant drawback of the approach is the high computational complexity and with that the lack of real-time guarantees for certain embedded systems integrated in aerial vehicles, for example the system integrated into the GHGV-2.
@@ 3 Problem Formulation
Figure 6: Rear view of the GHGV-2 showing available control effectors during endoatmospheric operations.
Figure 6 illustrates the available aerodynamic control surfaces and their corresponding deflections: upper left fin $u_{1}$, upper right fin $u_{2}$, lower left fin $u_{3}$, and lower right fin $u_{4}$ [Autenrieb2024]. This leads to the following control input vector:
@@
with $\mathbf{B}(\bm{x}(t),\bm{\sigma}(t))\in\mathbb{R}^{3\times 4}$. As it is a non-square matrix, it cannot be inverted directly. A major challenge in hypersonic flights, often neglected in the literature, is the dependence of actuator magnitude and rate limits not only on the instantaneous flight state $\bm{x}(t)$ but also on the operating point $\bm{\sigma}(t)$. During high-speed atmospheric operations, high dynamic pressure induces counterforces on actuators, limiting the achievable deflections and rates. Nonlinear aerodynamic effects further complicate matters, with asymmetric influences on upper and lower surfaces requiring distinct actuator constraints. Additionally, operating-point-dependent thermal loads must be considered to minimize heating, reduce the burden on the TPS, and lower the infrared signature of the vehicle during operations.
@@
These effects are illustrated in Figure 7. In this figure, the red arrow $\mathbf{V}$ represents the resulting velocity vector of the vehicle as it moves at high speed through the atmosphere. The blue arrows depict the airflow vector fields around the vehicle, shaping the aerodynamic forces and heating effects on the control surfaces. The control surface $i$, modeled as a flapping plate, is characterized by a deflection angle $u_{i}$ and a deflection rate $\dot{u}_{i}$. The red gradient along the control surface indicates aerodynamic heating, generated by the vehicle’s hypersonic speed, causing significant heat flux $\tau_{i}$ on the surface. These heating effects must be considered during operation to avoid material degradation and structural failure of the control surfaces and to minimize the infrared signature in defense scenarios. The green arrow $F_{q,i}$ represents the normal of the aerodynamic force acting on the surface, which is highly dependent on the deflection angle and the dynamic pressure $q=0.5\rho V^{2}$ acting on the control surface area $S_{u,i}$, with $\rho$ being the atmospheric density and $V$ the resulting velocity. The actuator force $\mathbf{F}_{A,i}$ is responsible for controlling the position of the surface, working against $\mathbf{F}_{q,i}$, especially under high dynamic pressure conditions. Under these conditions, the actuator’s ability to adjust the deflection angle and rate is sometimes restricted, leading to cases in which the deflection angle $u_{i}$ is limited such that the resulting dynamic pressure force $F_{q,i}$ exceeds the maximum force that the actuator can provide.
@@
with $\bm{u}_{s}(t)$ denoting a desired steady-state control input for a given virtual control command $\bm{\nu}(t)$. The vector $\bm{u}_{s}(t)$ represents a preferred steady-state solution of the algebraic allocation relation $\bm{\nu}(t)=\mathbf{B}(\mathbf{x}(t),\bm{\sigma}(t))\,\bm{u}(t)$ and serves as a reference point for the solution process. At this stage, $\bm{u}_{s}(t)$ is introduced as an auxiliary design quantity and is neither required to be unique nor enforced as a hard constraint. Its purpose is to encode designer preferences regarding steady-state actuator usage, for example with respect to load distribution or thermal considerations. The specific construction of $\bm{u}_{s}(t)$ and its integration within the subsequently proposed control allocation strategy are addressed in Section 5.2. The weighting matrix $\mathbf{W}_{m}$ penalizes deviations of the computed control input from the preferred steady-state solution, whereas $\mathbf{W}_{r}$ penalizes variations between consecutive control inputs $\bm{u}(t)$ and $\bm{u}(t-T)$, thereby promoting smooth actuator behavior. Design guidelines and problem-specific choices for the weighting matrices $\mathbf{W}_{m}$ and $\mathbf{W}_{r}$ for the control allocation problem of the GHGV-2 are provided in Section 5.1.
In summary, the problem addressed in this work is to determine, at each time instant, a control input $\bm{u}(t)$ for the system defined in Eqs. (1) and (2), in accordance with Eq. (31), such that the virtual control input demand $\mathbf{B}(\bm{x}(t),\bm{\sigma}(t))\,\bm{u}(t)=\bm{\nu}(t)$ is satisfied whenever the command lies within the state- and operating-point-dependent actuator limits. If the command is not feasible, the control input is determined so as to minimize the deviation between the requested and achievable virtual control input while respecting the input constraints defined in Eq. (30).
@@ 4 The Integrated Overall Control Allocation Concept
As discussed in Section 3, during atmospheric flight, HGVs face significant challenges in maintaining stability and controllability due to thermal loads and limited aerodynamic authority [Li_2017]. These constraints complicate the design of effective control strategies and require careful management of aerodynamic and thermal effects. Owing to this complexity, the nominal atmospheric flight trajectory of the HGV is typically computed offline prior to the mission. During this offline optimization process, an optimal control problem is solved to obtain a feasible reference trajectory that satisfies prescribed initial and terminal conditions as well as physical constraints [An_2020].
@@
The resulting optimized state reference trajectory is bounded by aerodynamic flight corridor constraints, as illustrated in Fig. 8 [Li_2017]. Exceeding the upper altitude bound leads to insufficient aerodynamic control authority due to low air density, whereas descending below the lower bound exposes the vehicle to excessive aerodynamic heating, potentially causing material degradation or structural failure. Based on the obtained state reference trajectory $\bm{x}_{r}(t)$, an offline baseline control input $\bm{u}_{r}(t)$ is computed using a computationally demanding but accurate QPCA method, as discussed in Section 2.2.3, for the problem defined in Eq. (31). This offline solution incorporates all known hard and soft actuator constraints and is assumed to represent an optimal solution for a disturbance-free environment. The resulting control inputs are stored in lookup tables and serve as baseline commands during flight. Under nominal conditions, the baseline control input $\bm{u}_{r}(t)$ is sufficient to maintain the vehicle on its reference trajectory. However, due to model uncertainties, unmodeled dynamics, and external disturbances inherent to hypersonic flight, additional corrective control action is required. To address this issue, a nonlinear feedback controller generates a corrective virtual control input $\Delta\bm{\nu}(t)$ to compensate for deviations from the nominal trajectory, further information about the controller are provided in [Autenrieb2024].
Assuming an affine structure of the control input, the total actuator command is expressed as
@@
Here, $\Delta\bm{u}(t)$ represents an incremental control input required to realize the corrective virtual control command $\Delta\bm{\nu}(t)$ that is generated by the nonlinear feedback controller. Although the assumed affine nature of the control input formally allows the decomposition and superposition of the baseline command and the incremental correction input, the proposed iterative online control allocation algorithm does not allocate $\Delta\bm{u}(t)$ independently. Instead, it searches for the full control input $\bm{u}(t)$ while using the offline-optimized allocation solution $\bm{u}_{r}(t)$ as a baseline. This design choice leverages the fact that $\bm{u}_{r}(t)$ already represents a physically meaningful and near-feasible control input for the nominal operating case, thereby providing a well-conditioned starting point for the online allocation task. By incorporating $\bm{u}_{r}(t)$ directly into the allocation process, the search for a feasible control input is biased toward solutions that remain close to the nominal baseline, which reduces the likelihood of rapid input changes and converging to poorly conditioned allocations. This is particularly important in scenarios with strongly asymmetric actuator effectiveness or state-dependent magnitude and rate constraints, where the admissible solution space may be highly restricted.
@@
The overall control allocation strategy is illustrated in Fig. 9. By combining an offline-optimized baseline solution with an online allocation mechanism operating on the full control input, the proposed architecture facilitates the real-time computation of feasible control commands in the presence of disturbances and uncertainties. The design and implementation of the iterative online control allocation algorithm, which constitutes a central contribution of this work, are detailed in the following section.
@@ 5 The Proposed Online Control Allocation Algorithm
In standard RPICA approaches, the problem of limited control input authority is addressed by iterating over the AMS until a feasible control input command is found [Virning_1994]. During the iteration steps, the algorithm typically employs the Moore–Penrose pseudoinverse. The main advantage of this approach is that a solution can often be obtained in a simple and computationally efficient manner, making it suitable for real-time implementation in many aerospace applications. While it is, in principle, possible to account for actuator rate information indirectly through equivalent magnitude constraints, such considerations are typically neglected in the majority of the practical RPICA implementations. Moreover, most RPICA formulations do not adequately represent asymmetric input or rate limitations, particularly when these constraints are intended to act as soft shaping mechanisms that influence the control distribution prior to constraint activation. For the class of hypersonic glide vehicles studied here, these aspects are particularly of high interest, rendering standard RPICA solutions suboptimal.
In the proposed iterative dynamic control allocation (IDCA) algorithm, a linear filter formulation, originally introduced in [Ola_2004], is employed. Unlike the Moore–Penrose inverse, this approach explicitly accounts for both the current control distribution and the allocation of the previous time step as soft constraints [Sadien_2020]. This implicitly incorporates actuator rates into the allocation problem. The method provides an explicit solution to the following weighted least-squares problem:
@@
where $\bm{\nu}(t)=\mathbf{B}\,\bm{u}_{r}(t)+\Delta\bm{\nu}(t)$ is the virtual control input to be generated. The desired steady-state solution $\bm{u}_{s}(t)$ is specified in Section 5.2. The algorithm is designed so that, if $\bm{u}_{s}(t)$ is feasible, the steady-state solution converges to $\bm{u}_{s}(t)$ or to a solution close to it. The weighting matrices $\mathbf{W}_{m}$ and $\mathbf{W}_{r}$ are symmetric and defined as functions of the state and control; their detailed design is discussed in Section 5.1. Since $\mathbf{W}_{m}$ and $\mathbf{W}_{r}$ are symmetric, the combined weighting
@@
The closed-form solution of Eq. (33), expressed in the form of Eq. (35), can be written as [Ola_2004]:
@@
The explicit solution of Eq. (36) is used within an iterative procedure to enforce actuator constraints. The algorithm is executed within a fixed controller cycle, so external variables such as $\bm{u}(t-T)$ and $\bm{x}(t)$ are treated as constant during the inner iterations. A closely related line of work was presented by Burken et al. [Burken_2001], who introduced a fixed-point, quadratic-programming-based iteration scheme for control allocation. Their method shares structural similarities with the iterative procedures considered here, in that both refine the solution by re-evaluating residual commands under actuator constraints. While the fixed-point scheme demonstrated promising performance for asymmetric re-entry vehicle configurations and provided a rigorous QP-based foundation for handling magnitude constraints, actuator rate limits were only mentioned as a possible extension and not explicitly included. By contrast, the algorithm presented in the following integrates rate-dependent limits directly into the allocation problem, both as explicit constraints and as soft penalties in the weighted least-squares cost. Moreover, the proposed formulation emphasizes practical implementation by illustrating how the weighting structure can incorporate operational considerations such as thermal load reduction and drag-sensitive allocation, and by providing empirical evidence in the simulation results that the method can robustly handle strongly time-varying, state- and operating-point-dependent actuator constraints.
@@
We define the iterative residual virtual control command as $\bm{\nu}^{(j)}$. In the first iteration step, the control input $\bm{u}^{(1)}$, corresponding to the initial residual $\bm{\nu}^{(0)}=\bm{\nu}(t)$, is computed using Eq. (36). If $\bm{u}^{(1)}$ satisfies all actuator magnitude and rate constraints and realizes the desired virtual control input $\bm{\nu}(t)$, it is accepted as the final solution. Otherwise, the iterative redistribution procedure is continued. The achieved saturated virtual control in iteration step $j$ is
@@
where $S(\cdot)$ is the state- and rate-dependent saturation function (cf. Eq. (29)), $\dot{\bm{u}}^{(j)}=\tfrac{1}{T}(\bm{u}^{(j)}-\mathbf{u}_{\tau})$ and $\mathbf{u}_{\tau}=\mathbf{u}(t-T)$ the control input of the last time step. If actuators saturate, the corresponding columns of $\mathbf{B}$ are set to zero, yielding the reduced matrix $\mathbf{B}_{s}$. Input limits are updated at each step to ensure feasibility:
@@
For completeness, alongside the mathematical steps and the process visualization in Figure 10, the algorithm is also summarized in pseudocode to complement the formulation and facilitate implementation.